If the system rotates perpendicular to the page at A and has a speed of 500rpm. Find a) rotational kinetic energy (KEr) b) angular momentum (L)
1
Expert's answer
2021-10-03T13:23:11-0400

Convert 500 rpm to rad/s:


"\omega=2\pi n=2\pi\frac{500}{60}=53.4\text{ rad/s}."

Rotational kinetic energy:


"E=\frac12 I\omega^2=1426I\text{ J}."

The angular momentum is


"L=I\omega=53.4I\text{ km}·\text{m}^2/\text{s}."

I is the moment of inertia of the system.